Researchers are interested in the association between depression and average hours of sleep received each night among a group of participants in an observational study aged 25-40. Researchers collected information about sleep habits by observing sleep journals collected from participants and compiling average sleep over the study period into a variable based on tertiles of the average sleep for the whole study population. The outcome was assessed by means of responses on the geriatric depression scale (GDS). Suppose the following table compiles the study results. Calculate the relative risks for depression by amount of sleep by comparing each of the upper tertiles to the 1st as the reference group. Depression (GDS>19) Depression (GDS≤19) Total 1st Tertile of Sleep 41 196 237 2nd Tertile of Sleep 56 181 237 3rd Tertile of Sleep 111 126 237 Total 208 503 711
Answer #1
Depression (GDS > 19) Depression (GDS ≤ 19) Total
1st Tertile of sleep 41 196 237
2nd Tertile of sleep 56 181 237
3rd Tertile of sleep 111 126 237
Total 208 503 711

Relative risk of depression(GDS > 19) in 2nd Tertile of sleep

= (56/237)/(41/237) = 56/41 = 1.366

Relative risk of depression (GDS > 19) in 3rd Tertile of sleep

= (111/237)/(41/237) = 111/41 = 2.707